# Successive Partial Disruptions with Orbital Precession in a White Dwarf–Black Hole System for Repeating GRB 250702B

### v1

@@ II.4 Jet Production and Luminosity
During each pericenter passage, a fraction of the WD mass is stripped and subsequently accreted onto the BH through an accretion disk. We assume that the accretion process launches a relativistic jet powered by the Blandford–Znajek (BZ) mechanism, which extracts rotational energy from the spinning BH via magnetic fields (Blandford and Znajek, 1977). Assuming that the jet has a half-opening angle $\theta_{\rm j}$ and that the emission is approximately uniform within the jet cone, the isotropic-equivalent gamma-ray luminosity is given by
$L_{\gamma,\mathrm{iso}}$ $\approx$ $\frac{\eta_{\rm BZ}\eta_{\rm rad}\,\Delta M\,c^{2}}{\theta_{\rm j}^{2}\,\delta t_{\mathrm{flare}}}$ (12)
$\sim$ $5\times 10^{51}~\eta_{\rm BZ}\left(\frac{\eta_{\rm rad}}{0.1}\right)\left(\frac{\delta t_{\rm flare}}{100~{\rm s}}\right)^{-1}\left(\frac{\Delta M}{2\times 10^{-2}~M_{\odot}}\right)$
$\times\left(\frac{\theta_{\rm j}}{0.1~{\rm rad}}\right)^{-2}~\rm{erg~s^{-1}},$
where $\eta_{\rm rad}$ is the radiative efficiency and $\eta_{\rm BZ}$ is the effective efficiency of the BZ jet, incorporating the effects of BH spin and accumulated magnetic flux. To reproduce the observed gamma-ray luminosity using the mass loss inferred from the observed timescales, a combined efficiency of $\eta_{\rm BZ}\eta_{\rm rad}\sim 0.1$ is required.
Such an efficiency implies that the BH is likely rapidly spinning ($a_{\rm spin}\sim 1$) and that the accretion disk is in a magnetically arrested disk state. Combining Eqs. (3), (9), and (11) under the condition $a_{\rm spin}\sim 1$ yields an upper limit on the BH mass of $M_{\rm BH}\lesssim 10^{6}\,M_{\odot}$, corresponding to a semi-major axis $a\lesssim 50\,R_{\odot}$.
@@ II.5 Precessing Orbits and Irregular Flare Intervals
Guided by the parameters inferred in Section II.3, we consider the BH mass $M_{\rm BH}\gtrsim 10^{5}\,M_{\odot}$, corresponding to $a\gtrsim 20\,R_{\odot}$ and $r_{\rm p}\lesssim 3\,r_{\rm g}$. In this regime, the WD orbit is well approximated by a bound Kerr geodesic, which undergoes relativistic precession (see Fig. 1(a) for an illustrative example).
To reproduce the fact that only a small number of gamma-ray flares are observed, the jet must lie outside the observer’s line of sight for most pericenter passages. This requires that the orbital precession angle per pericenter passage, $\theta_{\rm prec,orbit}\gtrsim 0.1~\mathrm{rad}$ ($\gtrsim 6^{\circ}$), exceed the jet half-opening angle adopted in Section II.4. Assuming a rapidly spinning BH ($a_{\rm spin}\sim 1$), solutions of the Kerr geodesic equations indicate that such precession is achieved for prograde orbits with the minimum polar angle of the Kerr geodesic measured from the BH equatorial plane $\theta_{\rm min}\gtrsim 0.12~\mathrm{rad}$ ($\theta_{\rm min}\gtrsim 7^{\circ}$). Although our geodesic calculations neglect mass loss, this has little impact at early times, when the WD orbit is largely unaffected by mass loss.
We assume that the accretion disk formed after each pericenter passage lies in the same plane as the WD orbit at pericenter, so that the disk angular-momentum vector is aligned with that of the WD. As a result, the disk (and the associated jet) is misaligned with respect to the BH spin axis by an angle $\theta_{\rm d}$. Such a misaligned disk experiences a Lense–Thirring torque, causing the disk and the associated jet to precess around the BH spin axis (e.g., Papaloizou and Lin, 1995; Fragile et al., 2007; Liska et al., 2018; Lu et al., 2024). Based on the mass-loss rate estimated in Section II.3, the accretion disk is expected to be in the super-Eddington regime, in which the misalignment between the disk–jet system and the BH spin axis can persist over timescales much longer than the Lense–Thirring precession timescale (e.g., Franchini et al., 2016; Zanazzi and Lai, 2019). The Lense–Thirring precession timescale is given by Lu et al. (2024) as $t_{\rm prec,LT}\gtrsim 100\,(r_{\rm in}/r_{\rm g})(M_{\rm BH}/10^{5},M_{\odot})\,{\rm s}$, which is comparable to or longer than the observed duration of a single gamma-ray flare ($\sim 100~{\rm s}$). This implies that the jet orientation can change by a non-negligible fraction of a precession cycle during a single flare. Therefore, the disk–jet system continues to undergo Lense–Thirring precession during an individual flare, with the jet sweeping out a cone of $\theta_{\rm d}$ around the BH spin axis.
@@ III Afterglow Emission
In the previous section, we consider that the WD can launch up to 40 jets over the entire sequence (hereafter, the maximum case). Most of these jets are initially viewed off-axis by the observer. As a result, the early-time emission is dominated by the four jets that happen to be launched nearly along the line of sight, while the remaining $\approx 36$ jets are initially invisible. As the jets decelerate, relativistic beaming weakens, allowing these off-axis jets to become visible to the observer (e.g., Granot et al., 2002). Their emission can therefore brighten at late times, particularly in the radio band (e.g., Lazzati et al., 2018; Sato et al., 2021, 2023). On the other hand, only four gamma-ray flares have been observed. This may indicate that only four jets were launched in total (hereafter referred to as the minimum case), possibly as a consequence of disk instabilities (Granot et al., 2025) and/or choked jet scenarios (Lu et al., 2024). Consequently, the maximum and minimum cases are expected to exhibit distinct late-time radio afterglow behaviors. In the following, we compare these two scenarios using their predicted afterglow emission.
@@
In the maximal case, we assume that jets are launched sequentially, one every hour, corresponding to the orbital period. For simplicity, the jet axis is assumed to remain fixed during the afterglow phase of each episode. The total afterglow emission is computed as the sum of the radiation from 40 jets with randomly assigned viewing angles between 0 and $\pi/2$. In the minimal case, jets are launched only at the times when gamma-ray flares are observed, and all jets are assumed to be viewed on-axis.
@@
We compute the multiwavelength afterglow emission from a single top-hat jet using the VegasAfterglow code (Wang et al., 2025). The jet is characterized by an initial half-opening angle $\theta_{0}$, bulk Lorentz factor $\Gamma_{0}$, and isotropic-equivalent kinetic energy $E_{\rm iso,K}$. It propagates into an external medium with a power-law density profile $n(r)=A_{*}(r/r_{0})^{-k}$, where $r_{0}=10^{17}\,{\rm cm}$, and produces synchrotron radiation from external shocks. We assume a power-law electron energy distribution with index $p$. The microphysical parameters $\epsilon_{e}$, $\epsilon_{B}$, and $f_{e}$ denote the fractions of internal energy in nonthermal electrons and magnetic fields, and the fraction of accelerated electrons, respectively, and are taken to be constant. The emission is calculated for a viewing angle $\theta_{\rm v,jet}$.
@@
We present our modeling of the multiwavelength afterglow emission in the X-ray (5 keV), near-infrared ($2\times 10^{14}$ Hz), and radio (1.3 GHz) bands, and compare the results with observations of GRB 250702B. The X-ray data are taken from Li et al. (2025) and O’Connor et al. (2025). We convert the observed 0.3–10 keV energy flux to a flux density at 5 keV, assuming a photon index of 1.8 during the slow-cooling phase. The near-infrared data are from Levan et al. (2025) and Carney et al. (2025). We adopt a V-band extinction of $A_{V}=4.2$ mag (Gompertz et al., 2025) and apply the corresponding extinction correction at $2\times 10^{14}$ Hz using a standard extinction law. The radio data are taken from Carney et al. (2025).
@@
In order to reproduce the observed multiwavelength afterglow light curves, we adopt the following model parameters: $\theta_{0}=0.1$ rad ($6^{\circ}$), $E_{\rm iso,K}=4\times 10^{54}$ erg, $\Gamma_{0}=100$, $k=1.0$, $A_{*}=0.1$, $p=2.6$, $\epsilon_{e}=10^{-2}$, $\epsilon_{B}=10^{-5}$, and $f_{e}=0.1$. Figure 2 presents the resulting multiwavelength emission. In both the maximum and minimum scenarios, the early afterglow is dominated by the same four jets viewed nearly on-axis, and therefore the early X-ray and near-infrared behavior is similar in both cases. The calculated X-ray light curve exhibits two pronounced flares between $\sim 4.2\times 10^{4}$ s and $\sim 1.2\times 10^{5}$ s, each corresponding to the transition of an on-axis jet from the coasting phase to the self-similar adiabatic expansion stage (Sari, 1997). The near-infrared light curve also shows two peaks, though with a different physical origin: each peak appears when the typical synchrotron frequency $\nu_{m}$ of the jet component crosses the observing band. Because jets are launched at different epochs, these peaks overlap in the observer frame and produce the sequence of rapidly rising and decaying flares in both X-ray and near-infrared bands.
@@
At later times, jets that were initially viewed off-axis in the maximum case decelerate sufficiently to become effectively on-axis (Granot et al., 2002). For $T\gtrsim 10^{6}\,{\rm s}$, the light curves of the maximum case therefore deviate from those of the four-jet configuration, as the off-axis jets begin to satisfy the condition $1/\Gamma\sim\theta_{\rm v,jet}-\theta_{\rm j}$ and enter the observer’s line of sight, enhancing the radio emission. This behavior is illustrated by the gray solid lines in Fig. 3, where each radio peak corresponds to the passage of the typical synchrotron frequency $\nu_{m}$ through 1.3 GHz. By $T\sim 10^{8}$ s, all jets have entered the Newtonian phase and can be regarded as effectively on-axis. The radio flux from the full jet population in the maximum case is then about an order of magnitude higher than that from the four-jet configuration. Because the microphysical parameters $A_{*}$, $\epsilon_{e}$, $\epsilon_{B}$, and $f_{e}$ are identical in both scenarios, the late-time radio emission reflects the total kinetic energy, $N_{\rm jet}E_{\rm iso,K}\theta_{0}^{2}/2$, where $N_{\rm jet}$ is the number of jets. At this epoch, $\nu_{m}$ lies below the radio band while the cooling frequency $\nu_{c}$ lies between the radio and near-infrared bands. The flux therefore scales as $F_{\nu}\propto E_{\rm iso,K}^{\frac{6+10p-k(5+p)}{4(5-k)}}A_{*}^{\frac{19-5p}{4(5-k)}}\epsilon_{B}^{\frac{p+1}{4}}\epsilon_{e}^{p-1}f_{e}^{2-p}\sim E_{\rm iso,K}^{1.5}A_{*}^{0.4}\epsilon_{B}^{0.9}\epsilon_{e}^{1.6}f_{e}^{-0.6}$. This scaling predicts the flux ratio of $(40/4)^{1.5}\sim 30$ between the maximum and minimum cases, consistent with our numerical results and yielding an order-of-magnitude difference in the radio flux at $T\sim 10^{8}$ s. At such late times, differences in jet launching epochs become negligible.
@@ IV Summary & Discussion
We have shown that the hierarchical timescales observed in the prompt emission of GRB 250702B can be explained by repeating partial tidal disruptions of the WD by the BH on a highly eccentric orbit ($e\approx 0.97$). For the rapidly spinning BH with $M_{\rm BH}\sim 10^{5}$$10^{6}\,M_{\odot}$ and $a_{\rm spin}\sim 1$, Kerr geodesics with the semi-major axis of $a\sim 20$$50\,R_{\odot}$ and the minimum polar angle measured from the BH equatorial plane of $\theta_{\rm min}\gtrsim 0.12~\mathrm{rad}$ ($\gtrsim 7^{\circ}$) undergo relativistic orbital precession, with the per-orbit precession angle exceeding the jet half-opening angle. This leads to changes in the jet direction between successive pericenter passages, explaining the irregular flare intervals. In our model, up to $\approx 40$ jets are launched during the sequence of partial disruptions, but only four are initially viewed on-axis and detected as prompt gamma-ray flares. The remaining $\approx 36$ jets are initially off-axis and invisible at early times, but their cumulative contribution enhances the late-time radio afterglow by about an order of magnitude, providing a testable signature of orbital precession in the WD–IMBH system.
The last three flares appear relatively clustered. Since the observer is located near the BH spin axis, mass loss during the late-stage passages may have gradually aligned the WD’s orbit with the BH’s equatorial plane. Once this configuration was established, orbital precession may have been suppressed, which could have kept the jet direction nearly constant from one pericenter passage to the next.
To account for the observed gamma-ray luminosity, a magnetic field strength of $B_{\rm H}\sim 1\times 10^{10}\,a_{\rm spin}(L_{\gamma,\mathrm{iso}}/5\times 10^{51}\,{\rm erg\,s^{-1}})^{\frac{1}{2}}(\theta_{\rm j}/0.1\,{\rm rad})(M_{\rm BH}/1\times 10^{5}\,M_{\odot})^{-1}(\kappa/0.06)^{-\frac{1}{2}}\,\mathrm{G}$ is required at the BH horizon, where $\kappa$ is a dimensionless coefficient appearing in the BZ formula (Tchekhovskoy et al., 2010, 2011). However, the magnetic field supplied by the WD material is expected to be much weaker, $B_{\rm adv}\sim 2\times 10^{5}~(B_{\rm WD}/10^{9}~{\rm G})(R_{\rm WD}/0.01~R_{\odot})^{2}(M_{\rm BH}/1\times 10^{5}~M_{\odot})^{-2}(\xi/0.1)\,\mathrm{G}$, where $B_{\rm WD}$ is the surface magnetic field of the WD and $\xi$ is the efficiency with which the WD magnetic field is transported to the BH horizon. This field strength is insufficient to power the jet. Therefore, an additional magnetic-field amplification mechanism, such as a dynamo process, would be necessary (e.g. Krolik and Piran, 2011; Tchekhovskoy et al., 2014).
@@
If our model is correct, this event implies the rapidly rotating IMBH ($M_{\rm BH}\sim 10^{5}\,M_{\odot}$, $a_{\rm spin}\sim 1$) hosting the WD on the highly eccentric orbit. Despite large uncertainties in their event rate (e.g., MacLeod et al., 2014; Ye et al., 2023), such systems are promising sources of low-frequency gravitational waves for space-based detectors such as the Laser Interferometer Space Antenna (e.g., Amaro-Seoane et al., 2023). Future multi-messenger observations may constrain the BH mass and spin, the orbital evolution, and enable tests of general relativity.
We thank Akihiro Inoue, Jim Fuller, Koki Kin, Yuki Kudoh, Tomoya Suzuguchi, and Kengo Tomida for valuable comments. This research was partially supported by JSPS KAKENHI Grant Nos. 25KJ0010 (YS), 24K17088 (TM), 22H00130, 23H04899, and 24K00668 ($K$).

### v2

@@ II.4 Jet Production and Luminosity
During each pericenter passage, a fraction of the WD mass is stripped. A portion of this material is accreted onto the BH through an accretion disk, while the rest is lost from the system (e.g., via outflows). We assume that the accretion process launches a relativistic jet powered by the Blandford–Znajek (BZ) mechanism, which extracts rotational energy from the spinning BH via magnetic fields (R. D. Blandford & R. L. Znajek, 1977). Assuming that the jet has a half-opening angle $\theta_{\rm j}$ and that the emission is approximately uniform within the jet cone, the isotropic-equivalent gamma-ray luminosity is given by
$L_{\gamma,\mathrm{iso,obs}}$ $\approx$ $\frac{2\eta_{\rm BZ}\eta_{\rm rad}(1-f)\,\Delta M\,c^{2}}{\theta_{\rm j}^{2}\,\delta t_{\mathrm{flare,obs}}}$ (12)
$\sim$ $5\times 10^{51}~\eta_{\rm BZ}\left(\frac{\eta_{\rm rad}}{0.1}\right)\left(\frac{\delta t_{\rm flare,obs}}{100~{\rm s}}\right)^{-1}\left(\frac{1-f}{0.5}\right)$
$\times\left(\frac{\Delta M}{2\times 10^{-2}~M_{\odot}}\right)\left(\frac{\theta_{\rm j}}{0.1~{\rm rad}}\right)^{-2}~\rm{erg~s^{-1}},$
where $f$ denotes the fraction of the stripped mass lost from the system (so that (1f)(1-f) is accreted onto the BH), $\eta_{\rm rad}$ is the radiative efficiency, and $\eta_{\rm BZ}$ is the effective efficiency of the BZ jet, incorporating the effects of BH spin and accumulated magnetic flux. To reproduce the observed gamma-ray luminosity using the mass loss inferred from the observed timescales, a combined efficiency of $\eta_{\rm BZ}\eta_{\rm rad}\sim 0.1$ is required.
Such an efficiency implies that the BH is likely rapidly spinning ($a_{\rm spin}\sim 1$) and that the accretion disk is in a magnetically arrested disk state. Combining Eqs. (3), (9), and (11) under the condition $a_{\rm spin}\sim 1$ yields an upper limit on the BH mass of $M_{\rm BH}\lesssim 10^{6}\,M_{\odot}$, corresponding to a semi-major axis $a\lesssim 60\,R_{\odot}$.
@@ II.5 Precessing Orbits and Irregular Flare Intervals
Guided by the parameters inferred in Section II.3, we consider the BH mass $M_{\rm BH}\gtrsim 10^{5}\,M_{\odot}$, corresponding to $a\gtrsim 30\,R_{\odot}$ and $r_{\rm p}\lesssim 3\,r_{\rm g}$. In this regime, the WD orbit is well approximated by a bound Kerr geodesic, which undergoes relativistic precession (see Fig. 1(a) for an illustrative example).
To reproduce the fact that only a small number of gamma-ray flares are observed, the jet must lie outside the observer’s line of sight for most pericenter passages. This requires that the orbital precession angle per pericenter passage, $\theta_{\rm prec,orbit}\gtrsim 0.1~\mathrm{rad}$ ($\gtrsim 6^{\circ}$), exceed the jet half-opening angle adopted in Section II.4. Assuming a rapidly spinning BH ($a_{\rm spin}\sim 1$), solutions of the Kerr geodesic equations indicate that such precession is achieved for prograde orbits with the minimum polar angle of the Kerr geodesic measured from the BH equatorial plane $\theta_{\rm min}\gtrsim 0.12~\mathrm{rad}$ ($\theta_{\rm min}\gtrsim 7^{\circ}$). Although our geodesic calculations neglect mass loss, this approximation is reasonable for the early passages considered, as the fractional mass loss per orbit is small.
We assume that the accretion disk formed after each pericenter passage lies in the same plane as the WD orbit at pericenter, so that the disk angular-momentum vector is aligned with that of the WD. As a result, the disk (and the associated jet) is misaligned with respect to the BH spin axis by an angle $\theta_{\rm d}$. Such a misaligned disk experiences a Lense–Thirring torque, causing the disk and the associated jet to precess around the BH spin axis (e.g., J. C. B. Papaloizou & D. N. C. Lin, 1995; P. C. Fragile et al., 2007; M. Liska et al., 2018; W. Lu et al., 2024; O. Teboul & B. D. Metzger, 2023). Based on the mass-loss rate estimated in Section II.3, the accretion disk is expected to be in the super-Eddington regime, in which the misalignment between the disk–jet system and the BH spin axis can persist over timescales much longer than the Lense–Thirring precession timescale (e.g., A. Franchini et al., 2016; J. J. Zanazzi & D. Lai, 2019). The Lense–Thirring precession timescale is given by W. Lu et al. (2024) as $t_{\rm prec,LT,obs}\gtrsim 200\,((1+z)/2)(r_{\rm in}/r_{\rm g})(M_{\rm BH}/10^{5}\,M_{\odot})\,{\rm s}$, which is longer than the observed duration of a single gamma-ray flare ($\sim 100~{\rm s}$). This implies that the jet orientation can change by a non-negligible fraction of a precession cycle during a single flare. Therefore, the disk–jet system continues to undergo Lense–Thirring precession during an individual flare, with the jet sweeping out a cone of $\theta_{\rm d}$ around the BH spin axis. As a consequence, the observed flare duration may depend on the instantaneous viewing geometry and precession phase, potentially leading to a moderate dispersion in flare lengths. Although the flares are typically of order $\sim 100$ s, the observed spread in durations is almost consistent with this picture.
@@ III Afterglow Emission
In the previous section, we consider that the WD can launch up to 40 jets over the entire sequence (hereafter, the maximum case). Most of these jets are initially viewed off-axis by the observer. As a result, the early-time emission is dominated by the four jets that happen to be launched nearly along the line of sight, while the remaining $\approx 36$ jets are initially invisible. As the jets decelerate, relativistic beaming weakens, allowing these off-axis jets to become visible to the observer (e.g., J. Granot et al., 2002). Their emission can therefore brighten at late times, particularly in the radio band (e.g., D. Lazzati et al., 2018; Y. Sato et al., 2021, 2023). On the other hand, only four gamma-ray flares have been observed. This may indicate that only four jets were launched in total (hereafter referred to as the minimum case), possibly as a consequence of disk instabilities (J. Granot et al., 2025) and/or choked jet scenarios (W. Lu et al., 2024). Consequently, the maximum and minimum cases are expected to exhibit distinct late-time radio afterglow behaviors. In the following, we compare these two scenarios using their predicted afterglow emission.
@@
In the maximal case, we assume that jets are launched sequentially, one every hour, corresponding to the orbital period. For simplicity, the jet axis is assumed to remain fixed during the afterglow phase of each episode. The total afterglow emission is computed as the sum of the radiation from 40 jets with randomly assigned viewing angles between 0 and $\pi/2$. Geodesic calculations show that for $\theta_{\rm min}\sim\pi/2$, the jet becomes nearly perpendicular to the BH spin axis, allowing viewing angles up to $\pi/2$ for an observer located near the spin axis. In the minimal case, jets are launched only at the times when gamma-ray flares are observed, and all jets are assumed to be viewed on-axis.
@@
We compute the multiwavelength afterglow emission from a single top-hat jet using the VegasAfterglow code (Y. Wang et al., 2025). The jet is characterized by an initial half-opening angle $\theta_{j}$, bulk Lorentz factor $\Gamma_{0}$, and isotropic-equivalent kinetic energy $E_{\rm iso,K}$. It propagates into an external medium with a power-law density profile $n(r)=A_{*}(r/r_{0})^{-k}$, where $r_{0}=10^{17}\,{\rm cm}$, and produces synchrotron radiation from external shocks. We assume a power-law electron energy distribution with index $p$. The microphysical parameters $\epsilon_{e}$, $\epsilon_{B}$, and $f_{e}$ denote the fractions of internal energy in nonthermal electrons and magnetic fields, and the fraction of accelerated electrons, respectively, and are taken to be constant. The emission is calculated for a viewing angle $\theta_{\rm v,jet}$.
@@
We present our modeling of the multiwavelength afterglow emission in the X-ray (5 keV), near-infrared ($2\times 10^{14}$ Hz), and radio (1.3 GHz) bands, and compare the results with observations of GRB 250702B. The X-ray data are taken from D. Y. Li et al. (2025) and B. O’Connor et al. (2025). We convert the observed 0.3–10 keV energy flux to a flux density at 5 keV, assuming a photon index of 1.8 during the slow-cooling phase. The near-infrared data are from A. J. Levan et al. (2025) and J. Carney et al. (2025). We adopt a V-band extinction of $A_{V}=4.2$ mag (B. P. Gompertz et al., 2025) and apply the corresponding extinction correction at $2\times 10^{14}$ Hz using a standard extinction law. The radio data are taken from J. Carney et al. (2025).
In order to reproduce the observed multiwavelength afterglow light curves, we adopt the following model parameters: $\theta_{j}=0.1$ rad ($6^{\circ}$), $E_{\rm iso,K}=4\times 10^{54}$ erg, $\Gamma_{0}=100$, $k=1.0$, $A_{*}=0.1$, $p=2.6$, $\epsilon_{e}=10^{-2}$, $\epsilon_{B}=10^{-5}$, and $f_{e}=0.1$. For a jet half-opening angle of $\theta_{j}\sim 0.1$ rad, the collimation-corrected kinetic energy is $E_{\rm jet,K}\sim E_{\rm iso,K}\theta_{0}^{2}/2\sim 2\times 10^{52}$ erg per episode, which is comparable to the energy available $\eta_{\rm BZ}\Delta Mc^{2}$.
@@
Figure 2 presents the resulting multiwavelength emission. In both the maximum and minimum scenarios, the early afterglow is dominated by the same four jets viewed nearly on-axis, and therefore the early X-ray and near-infrared behavior is similar in both cases. The calculated X-ray light curve exhibits two pronounced flares between $\sim 4.2\times 10^{4}$ s and $\sim 1.2\times 10^{5}$ s, each corresponding to the transition of an on-axis jet from the coasting phase to the self-similar adiabatic expansion stage (R. Sari, 1997). The near-infrared light curve also shows two peaks, though with a different physical origin: each peak appears when the typical synchrotron frequency $\nu_{m}$ of the jet component crosses the observing band. Because jets are launched at different epochs, these peaks overlap in the observer frame and produce the sequence of rapidly rising and decaying flares in both X-ray and near-infrared bands.
@@
At later times, jets that were initially viewed off-axis in the maximum case decelerate sufficiently to become effectively on-axis (J. Granot et al., 2002). For $T\gtrsim 10^{6}\,{\rm s}$, the light curves of the maximum case therefore deviate from those of the four-jet configuration, as the off-axis jets begin to satisfy the condition $1/\Gamma\sim\theta_{\rm v,jet}-\theta_{\rm j}$ and enter the observer’s line of sight, enhancing the radio emission. This behavior is illustrated by the gray solid lines in Fig. 3, where each radio peak corresponds to the passage of the typical synchrotron frequency $\nu_{m}$ through 1.3 GHz. By $T\sim 10^{8}$ s, all jets have entered the Newtonian phase and can be regarded as effectively on-axis. The radio flux from the full jet population in the maximum case is then about an order of magnitude higher than that from the four-jet configuration. Because the microphysical parameters $A_{*}$, $\epsilon_{e}$, $\epsilon_{B}$, and $f_{e}$ are identical in both scenarios, the late-time radio emission reflects the total kinetic energy, $N_{\rm jet}E_{\rm iso,K}\theta_{j}^{2}/2$, where $N_{\rm jet}$ is the number of jets. At this epoch, $\nu_{m}$ lies below the radio band while the cooling frequency $\nu_{c}$ lies between the radio and near-infrared bands. The flux therefore scales as $F_{\nu}\propto E_{\rm iso,K}^{\frac{6+10p-k(5+p)}{4(5-k)}}A_{*}^{\frac{19-5p}{4(5-k)}}\epsilon_{B}^{\frac{p+1}{4}}\epsilon_{e}^{p-1}f_{e}^{2-p}\sim E_{\rm iso,K}^{1.5}A_{*}^{0.4}\epsilon_{B}^{0.9}\epsilon_{e}^{1.6}f_{e}^{-0.6}$. This scaling predicts the flux ratio of $(40/4)^{1.5}\sim 30$ between the maximum and minimum cases, consistent with our numerical results and yielding an order-of-magnitude difference in the radio flux at $T\sim 10^{8}$ s. At such late times, differences in jet launching epochs become negligible.
@@ IV Summary & Discussion
We have shown that the hierarchical timescales observed in the prompt emission of GRB 250702B can be explained by repeating partial tidal disruptions of the WD by the BH on a highly eccentric orbit ($e\approx 0.97$). For the rapidly spinning BH with $M_{\rm BH}\sim 10^{5}$$10^{6}\,M_{\odot}$ and $a_{\rm spin}\sim 1$, Kerr geodesics with the semi-major axis of $a\sim 30$$60\,R_{\odot}$ and the minimum polar angle measured from the BH equatorial plane of $\theta_{\rm min}\gtrsim 0.12~\mathrm{rad}$ ($\gtrsim 7^{\circ}$) undergo relativistic orbital precession, with the per-orbit precession angle exceeding the jet half-opening angle. This leads to changes in the jet direction between successive pericenter passages, explaining the irregular flare intervals. In our model, up to $\approx 40$ jets are launched during the sequence of partial disruptions, but only four are initially viewed on-axis and detected as prompt gamma-ray flares. The remaining $\approx 36$ jets are initially off-axis and invisible at early times, but their cumulative contribution enhances the late-time radio afterglow by about an order of magnitude, providing a testable signature of orbital precession in the WD–IMBH system.
The last three flares appear relatively clustered. Since the observer is located near the BH spin axis, one possible interpretation is that the minimum polar angle, $\theta_{\min}$, could have evolved on the secular timescale, bringing the WD orbit closer to the BH equatorial plane. If so, this change in the orbital configuration may have led to a more stable jet direction, keeping the jet orientation nearly constant from one pericenter passage to the next.
To account for the observed gamma-ray luminosity, a magnetic field strength of $B_{\rm H}\sim 2\times 10^{10}\,((1+z)/2)a_{\rm spin}(L_{\gamma,\mathrm{iso,obs}}/5\times 10^{51}\,{\rm erg\,s^{-1}})^{\frac{1}{2}}(\theta_{\rm j}/0.1\,{\rm rad})(M_{\rm BH}/1\times 10^{5}\,M_{\odot})^{-1}(\kappa/0.06)^{-\frac{1}{2}}\,\mathrm{G}$ is required at the BH horizon, where $\kappa$ is a dimensionless coefficient appearing in the BZ formula (A. Tchekhovskoy et al., 2010, 2011). However, the magnetic field supplied by the WD material is expected to be much weaker, $B_{\rm adv}\sim 2\times 10^{5}~(B_{\rm WD}/10^{9}~{\rm G})(R_{\rm WD}/0.01~R_{\odot})^{2}(M_{\rm BH}/1\times 10^{5}~M_{\odot})^{-2}(\xi/0.1)\,\mathrm{G}$, where $B_{\rm WD}$ is the surface magnetic field of the WD and $\xi$ is the efficiency with which the WD magnetic field is transported to the BH horizon. This field strength is insufficient to power the jet. Therefore, an additional magnetic-field amplification mechanism, such as a dynamo process, would be necessary (e.g. J. H. Krolik & T. Piran, 2011; A. Tchekhovskoy et al., 2014).
@@
If our model is correct, this event implies a rapidly rotating IMBH ($M_{\rm BH}\sim 10^{5}\,M_{\odot}$, $a_{\rm spin}\sim 1$) hosting a WD on a highly eccentric orbit. Despite large uncertainties in their event rate (e.g., M. MacLeod et al., 2014; C. S. Ye et al., 2023), such systems are promising sources of low-frequency gravitational waves for space-based detectors such as the Laser Interferometer Space Antenna (LISA) (e.g., P. Amaro-Seoane et al., 2023). For IMBH mass, the characteristic orbital frequencies fall in the mHz band, close to the peak sensitivity of LISA. Recent studies have highlighted BH–WD binaries (e.g., Q.-Q. Ma et al., 2025; Z. Xuan et al., 2025), as promising multimessenger sources in the mHz band. Future multi-messenger observations may constrain the BH mass and spin, the orbital evolution, and enable tests of general relativity.
We thank Akihiro Inoue, Jim Fuller, Jonathan Granot, Koki Kin, Yuki Kudoh, Tomoya Suzuguchi, and Kengo Tomida for valuable comments. We also thank the anonymous referee and the editor for their helpful comments that improved the paper. The authors are grateful to the Yukawa Institute for Theoretical Physics at Kyoto University. This work was completed in part during the YITP long-term workshop YITP-T-26-02, “Multi-Messenger Astrophysics in the Dynamic Universe.” This research was partially supported by JSPS KAKENHI Grant Nos. 25KJ0010 (YS), 24K17088 (TM), 22H00130, 23H04899, and 24K00668 ($K$).